# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ 5 Constraining the partitioning of tidal heating between the shell and the core
The total tidal energy dissipated in Enceladus can be computed by measuring the amplitude and phase lag of $k_{2}$, independent of the choice of rheological model (see Equation (5)). We can compute the tidal heating in any specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body. Here, we use equation 37 in Tobie et al. (2005) to derive the contribution of the core and the shell to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (12)
where $H_{tidal}$ is the radial distribution of the dissipation rate per unit volume, $r$ is the radius from the center, $H_{\mu}$ is the sensitivity kernel for shear dissipation given in Equation (33) in Tobie et al. (2005), and $\Im{(\mu(r))}$ is the imaginary part of the complex shear modulus which is assumed constant at each layer. By integrating this expression from the center of the satellite to the core-ocean boundary, we obtain:
@@
The integral can be written in terms of the imaginary part of the piecewise tidal Love number of the core following Equation 35 in Tobie et al. (2005):
[…]
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational Love number associated with the core. Thus, if we compute the contribution of each of the layers to $|k_{2}|/Q$ or identically to $\Im(k_{2})$ of the body, that is, evaluating at each interface, we can compute the associated tidal heating for the part of the body that is interior to the interface. For a homogeneous core, $\hat{k}_{2}^{core}$ can be computed from the deformation at the top of the core:
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (15)
Tidal heating in the ice shell is then computed by subtracting the contribution of the core from the total energy dissipation:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (16)
Equations (13) and (16) provide a method to calculate the partition of the global tidal heating for a given interior structure model. These equations require a priori assumptions or independent constraints on the rheological and structural properties to compute the shear dissipation in the core. Therefore, here we derive the partitioning of tidal heating on the basis of the observable quantities, i.e., the tidal Love numbers. We extend the formulation provided by Beuthe (2019) which uses membrane mechanics to obtain the contribution of each layer of the body to the imaginary part of the tidal Love number. The shell’s contribution can be written as (Beuthe, 2019):
where, following the terminology used in Beuthe (2019), $\xi_{2}=3\rho_{ocean}/5\rho_{E}$, $\rho_{E}$ is the density of the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, and $k_{2}^{0}$ is the fluid-crust gravitational Love number (assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). In Appendix B, we show that the tidal dissipation occurring in the shell can be expressed in terms of the tidal Love numbers and density of the layers:
Here, $\delta\rho$ is the density jump from the ocean to the crust and $\epsilon$ is the ratio of the thickness of the crust to Enceladus’s radius. Equation (18) expresses the tidal heating in the shell using measurements of the real and imaginary parts of the Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (17). Tidal heating in the core can be accordingly computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
6 Road-map to Constrain the interior properties with geodetic observations
Despite numerous studies based on the observations made by Cassini, multiple interior structure properties of Enceladus are not well constrained. This uncertainty is due in part to the limitation of the available observations and in part to differences in the interpretation of the data for the interior structure and the underlying modeling assumptions (e.g., Howett et al., 2011; Spencer et al., 2013; Thomas et al., 2016; Park et al., 2024; Schenk & McKinnon, 2024). Here, we explore the ranges of possible interior parameters that are consistent with the currently available observations. Based on these constraints, we predict ranges of the currently unavailable geodetic observables, and investigate the possibility of discriminating between the two main tidal heating scenarios from future measurements. We use the estimates on the observables as a basis for exploring the interior from future measurements.
6.1 Constraining the interior and discriminating between heat generation scenarios
Currently available observations that can be used to understand Enceladus’s interior structure and heat production include shape (e.g., Park et al., 2024; Schenk & McKinnon, 2024), gravity up to degree-3 (Iess et al., 2014; Park et al., 2024), physical libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024), and heat flux (Howett et al., 2011; Spencer et al., 2013) as summarized in Table 1.
Figure 4: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude spanned by estimates in Thomas et al. (2016), Nadezhdina et al. (2016) and Park et al. (2024). Other observables in this plot are free to vary. In the 2D Probability Density Functions (PDFs), black (yellow) color indicates low (high) probability. The complex patterns observed indicating multiple peaks in the PDFs of $k_{2}$, $h_{2}$, and $l_{2}$ and phase lags reflect the different regimes of dissipation consistent with the observations (see Figure 2), resulting in the nonlinearity of the forwards model. Phase lags of $k_{2}$, $h_{2}$, and $l_{2}$ are in degrees and amplitude of the libration is in meters. $J_{3}$ is shown divided by 106.
We start with the most conservatively chosen set of observational constraints to show the effect of incorporating various sets of observations on the measurement requirements for future exploration. We assume spherical symmetry and consider a set of interior parameters, i.e., thickness, rigidity, effective viscosity, and density of the layers, and geodetic observable quantities, i.e., tidal Love numbers, phase lag angles, libration, moment of inertia, and $l=3$ zonal gravity coefficient, hereafter referred to as observables. We conduct Monte Carlo sampling based on the prior ranges of the model parameters presented in Table 2 (see Appendix C). We present the correlation between the observables in Figure 4. We use the total density, the physical libration amplitude in a range covering the estimates by Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024), and the estimated heat loss rate as the only constraints. To demonstrate the widest potentially plausible parameter space, we have also considered a wide range of published estimates of the heat loss rate suggested in Howett et al. (2011), Hemingway & Mittal (2019), and Park et al. (2024) (15 GW$<\dot{E}<$40 GW). The Love numbers, phase lags, moment of inertia, and degree-3 static gravity coefficient vary freely as long as the model satisfies the constraints of mass, heat loss, and libration; that is, we do not impose any constraints on the estimated moment of inertia and $J_{3}$ to avoid any possible inconsistency of our models with the underlying assumptions used in deriving these estimates.
@@ 6.1 Constraining the interior and discriminating between heat generation scenarios
Consistent with Section 3, the patterns of the posterior Probability Density Functions (PDFs) of the amplitude and phase lag of $k_{2}$ largely differ from the PDFs of the amplitudes and phases of $h_{2}$ and $l_{2}$ (Figure 4). PDFs of $k_{2}$ and $\epsilon_{k_{2}}$ have several peaks that are associated with tidal activity dominated in the core versus in the shell. A different pattern of distribution and peaks associated with different regimes of tidal activity is observed for $h_{2}$ and $l_{2}$, and their phase lags, since these quantities are largely insensitive to tidal activity in the core. The posterior PDFs show that while the $h_{2}$ and $l_{2}$ and their phase lags (Figure 4) are strongly correlated with each other (dominated by the response of the shell), their correlation with the amplitude and phase lag of $k_{2}$ depends on where the tidal dissipation dominantly occurs.
@@
We also obtain the ranges of the interior model parameters and the correlation between them associated with the observables (Appendix Figure 18). Previously, we ruled out mean crustal thicknesses around 5–10 km for Enceladus, as they would result in amplitudes of libration that are too large due to resonance in the shell (see Section 4). Ice shell thicknesses $\approx$ 1–5 km would be compatible with the estimates of tidal heating and libration only if the mean effective viscosity of the shell is unrealistically low. Therefore, we conclude that the mean thickness of Enceladus’s shell is larger than 10 km. For completeness, we included very low effective viscosities for the shell and the core, which imply deformations in the viscous-dominated regimes. As a result, the PDFs of the viscosities of the shell and core indicate multiple peaks representative of tidal dissipation dominated by viscous deformation as well as the anelastic deformation (See also Figure 1). From here on out, we consider viscosities of the ice shell and the rocky core that are close to or greater than their nominal relaxation viscosities ($\eta_{\rm shell}~{}>~{}10^{12}$ Pa.s and $\eta_{\rm core}~{}>~{}10^{11}$ Pa.s), i.e., the ice shell and the rocky core deform in the anelasticity-dominated deformation regime and not in the viscous-dominated deformation regime.
@@
We focus on analysis of discriminating between two sets of geodetic observables and interior parameters that are associated with each of the two scenarios (Figures 5, and Figures 6). We compute the tidal heating in the shell using Equation (16). We differentiate between the core-dominated and shell-dominated tidal dissipation scenarios by defining a threshold whereby at least 90% of the total heat is produced in the shell or the core. The PDFs of $k_{2}$ and $\epsilon_{k_{2}}$ for the two tidal dissipation scenarios largely overlap (Figures 5 and 6), whereas the PDFs of $h_{2}$, $\epsilon_{h_{2}}$, $l_{2}$, and $\epsilon_{l_{2}}$ in the case of shell-dominated tidal dissipation (Figure 5) differ significantly from those in the case of core-dominated tidal dissipation (Figure 6). In an inverse problem context, the degeneracy caused by the two different tidal dissipation scenarios reflected in the tidal Love numbers, as discussed earlier, leads to the non-uniqueness of the plausible interpretation of $k_{2}$ and $\epsilon_{k_{2}}$.
@@
We obtain tight constraints on the viscosity of the shell in the case where we impose the constraint that the tidal activity is dominated in the shell, and on the viscosity of the core in the case where a core-dominated tidal dissipation occurs (Figure 7). The rheological properties of the region in which tidal dissipation is not focused remain largely unconstrained. Using this set of observables, we find that the thickness of the ice shell, mostly constrained by the libration amplitude, is in the range of 15–40 km. Since we have conservatively chosen broad ranges for observables in this analysis, we have not obtained tight constraints on most of the other model parameters in this analysis. The PDFs and correlations between the interior structure parameters are presented in Appendix Figures 19 and 20.
@@
Figure 5: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and the libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024). Here, we assume that the dissipation is dominantly occurring in the ice shell. Unit conventions are the same as in Figure 4.
Figure 6: Exploration of the potentially observable geodetic quantities consistent with mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024). Here, we assume that tidal dissipation occurs dominantly in the core. Unit conventions are the same as in Figure 4.
@@
Figure 7: PDFs of ice shell and core viscosities for each of the two cases of tidal dissipation occurring dominantly in the shell (purple) and in the core (yellow), corresponding to Figure 5 and Figure 6, respectively.
@@
We compare the PDFs of $h_{2}$ and $l_{2}$ and their phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) associated with the two heat generation scenarios in Figure 8. The 2-$\sigma$ standard deviations are indicated in this figure. The separation between the two intervals associated with the 2-$\sigma$ uncertainties corresponding to the two PDFs determines the measurement precision required to discriminate between the two dissipation regimes. In other words, if we measure $h_{2}$ with a precision equal to the separation between the 2-$\sigma$ confidence intervals (or 1-$\sigma$, not shown in the figure) of the two PDFs, we obtain a confidence $\sim$95(or 68% if 1-$\sigma$ intervals are used) in identifying the layer that dominates the dissipation budget. Due to the overlap between the PDFs of $h_{2}$ and $l_{2}$ associated with the two tidal dissipation scenarios using this set of observations, distinguishing between the two cases is challenging. This large overlap is due to the large uncertainties in the observables considered in this case, and taking into account only the currently available constraints. The analysis implies that measurement of the maximum vertical tidal deformations with a maximum uncertainty of 62 cm can distinguish between the dominant region of tidal activity within the 1-$\sigma$ confidence level. Measurement of radial deformations with a maximum uncertainty of 56 cm can indicate the more “likely” region of heat generation, but it cannot definitively discriminate between the two scenarios (This value is obtained by the separation associated with the 2-$\sigma$ uncertainties between the two regimes which overlap in this case). Alternatively, measurement of the tangential displacements within an uncertainty of 28 cm is required for 1-$\sigma$ confidence. The phase lags associated with the radial deformations ($\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$) indicate a better separation between the PDFs associated with the two regimes of tidal activity compared to the deformations and $h_{2}$. We find that measuring the phase lag angles of the tidal deformations with an accuracy of 11 is required for 1-$\sigma$ confidence, and 4 is required for 2-$\sigma$ confidence in determining the dominant region of tidal dissipation.
Figure 8: Vertical and horizontal displacement Love numbers ($|h_{2}|$, $|l_{2}|$) and phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) are presented for the core-dominated and shell-dominated tidal dissipation. The heat flux is assumed to be $\rm 15~{}GW<\dot{E}<40~{}GW$, which includes the lowest and highest ranges estimated in Hemingway et al. (2018) and Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma_{I}<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). The top axis in the plots for $h_{2}$ and $l_{2}$ indicate the associated maximum radial and maximum easterly displacements corresponding to the values of the Love number in the bottom axis, respectively.
@@
Next, we focus on a narrower range of observational constraints using exclusively the values and uncertainties in Park et al. (2024) for $\gamma_{I}$, $J_{3}$, moment of inertia, mean density, and the total heat loss rate (see Table 1) to constrain the interior structure. We use 3-$\sigma$ uncertainties on the observables. We present the correlation between the observables and the model parameters obtained based on this analysis for a case that incorporates both tidal dissipation scenarios in Figures 9a and 9b. The obtained preferred values of the observables and uncertainties based on Park et al. (2024), including the tidal Love numbers, phase lags, moment of inertia, libration amplitude, and $J_{3}$ are given in Table 3. We also quantify the values for the observables that present the highest likelihood for each of the two heat generation scenarios. We will use these synthetic observations to perform inversions using plausible uncertainties expected from future measurements. Using the narrower observational constraints based on Park et al. (2024) further demonstrate that while $k_{2}$ and $\epsilon_{k_{2}}$ show a normal distribution with a single peak, $h_{2}$, $l_{2}$, $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ show two peaks associated with the two tidal dissipation scenarios. We note that satisfying all the measurements provided in Park et al. (2024) is possible when using 3-$\sigma$ uncertainties on the observables. Specifically, we find that the mean value of $J_{3}$ can be matched with the libration amplitude when considering 3-$\sigma$ uncertainty on the measured $J_{3}$. In contrast, Genova et al. (2024) use only the gravity measurements without exploiting the libration measurement as a constraint, which is expected to somewhat alter the results. Our results can also be compared to those presented by Genova et al. (2024) and Ermakov et al. (2021) where only tidal heating in the shell is considered. We also obtain considerably improved constraints on the interior parameters (Figure 9b) compared to the case where we use broader uncertainties on the observations. However, several model parameters remain weakly constrained with the currently available observations. The ranges of the obtained plausible interior parameters are listed in Table 4. See Appendix Figures 21 and 22 for the 2D PDFs of the model parameters associated with each of the cases.
@@
Table 3: Favored values and uncertainties (3-$\sigma$) of observables calculated using MCMC sampling based on the available inferred observables given by Park et al. (2024). Also, expected uncertainties from future explorations are provided (see Section 6.4) which are used in the inversion in Section 6.2. The values in the parentheses in the expected measurements of the tidal Love numbers and phase lag corresponds to a scenario with lower accuracy, as discussed in the text.
@@
$\mu_{\rm shell}$ GPa $2-4$ $2-4$
$\rho_{\rm shell}$ kg/m3 $700-950$ $700-950$
$\rho_{\rm ocean}$ kg/m3 $920-1250$ $920-1250$
$R_{\rm core}$ km $183-213$ $183-213$
$\mu_{\rm core}$ GPa $4-70$ $5-80$
$\rho_{\rm core}$ kg/m3 $2160-2400$ $2160-2400$

### v3

@@ 5 Constraining the partitioning of tidal heating between the shell and the core
The total tidal energy dissipated in Enceladus can be computed by measuring the amplitude and phase lag of $k_{2}$, independent of the choice of rheological model (see Equation (5)). We can compute the tidal heating in any specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body. Here, we use equations (33), (35), and (37) in Tobie et al. (2005) to derive the contribution of the core and the shell to the global tidal heating:
$\dot{E}_{core}=-\frac{21\pi}{2}\omega^{5}e^{2}R_{E}^{4}\frac{R_{core}}{\pi G}\Im{(\hat{k}_{2}^{core}}).$ (12)
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational Love number associated with the core. Thus, if we compute the contribution of each of the layers to $|k_{2}|/Q$ or identically to $\Im(k_{2})$ of the body, that is, evaluating at each interface, we can compute the associated tidal heating for the part of the body that is interior to the interface. For a homogeneous core, $\hat{k}_{2}^{core}$ can be computed from the deformation at the top of the core ($\Delta R_{core}$):
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}},$ (13)
Details of this derivation are presented in Appendix C. The tidal heating in the ice shell is then calculated by subtracting the contribution of the core from the total energy dissipated:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (14)
Tidal dissipation in the ocean is assumed negligible (Hay & Matsuyama, 2017; Matsuyama et al., 2018)
Equations (12) and (14) provide a method to calculate the partitioning of the global tidal heating for a given interior structure model. These equations require a priori assumptions or independent constraints on the rheological and structural properties to compute the shear dissipation in the core. Therefore, here we derive the partitioning of tidal heating on the basis of the observable quantities, i.e., the tidal Love numbers. We extend the formulation provided by Beuthe (2019) which uses membrane mechanics to obtain the contribution of each layer of the body to the imaginary part of the tidal Love number. The shell’s contribution can be written as (Beuthe, 2019):
where, following the terminology used in Beuthe (2019), $\xi_{2}=3\rho_{ocean}/5\rho_{E}$, $\rho_{E}$ is the density of the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, and $k_{2}^{0}$ is the fluid-crust gravitational Love number (assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). In Appendix C, we show that the tidal dissipation occurring in the shell can be expressed in terms of the tidal Love numbers and density of the layers:
Here, $\delta\rho$ is the density difference between the ocean and the crust. $\epsilon$ is the ratio of the thickness of the crust to Enceladus’s radius. Equation (16) expresses the tidal heating in the shell using measurements of the real and imaginary parts of the Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers required in Equation (15). Based on Equation (16), the uncertainty in the magnitude of tidal heating occurring in the shell can be computed from the uncertainties on the density structure and the complex Love numbers. The tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
6 Road-map to Constrain the interior properties
Despite numerous studies based on Cassini observations, numerous interior structure properties of Enceladus are poorly constrained. This uncertainty is due in part to the limitation of the available observations and in part to differences in the interpretation of the data for the interior structure and the underlying modeling assumptions (e.g., Howett et al., 2011; Spencer et al., 2013; Thomas et al., 2016; Park et al., 2024; Schenk & McKinnon, 2024). Here, we explore the ranges of possible interior parameters that are consistent with the currently available observations. Based on these constraints, we predict ranges of the currently unavailable geodetic observables, and investigate the possibility of discriminating between the two main tidal heating scenarios from future measurements. We use the estimates of the observables as a basis for exploring the interior from future measurements.
6.1 Constraining the Interior and Discriminating Between Dissipation Scenarios
Currently available observations that can be used to understand Enceladus’s interior structure and heat production include shape (e.g., Park et al., 2024; Schenk & McKinnon, 2024), gravity up to degree-3 (Iess et al., 2014; Park et al., 2024), physical libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024), and heat flux (Howett et al., 2011; Spencer et al., 2013) as summarized in Table 1. We highlight that in this paper, in the absence of more evidence, we assume that the total heat loss is equal to the tidally generated heat computed from Equation (5); in other words, we assume that Enceladus is currently in a thermal steady state. Moreover, we also note that the rates of heat loss considered here, except by Howett et al. (2011), are not based on direct measurements, but are estimated from the shell thickness and assumptions on its thermal conductivity. The rate of heat loss is largely used to determine the tidal phase lag. To avoid any self-consistency issue, we ensure that the obtained a posteriori ice shell thicknesses are within the ranges used to determine the adopted rates of heat loss. We further discuss the rate of heat loss in Enceladus in Section 7.
Figure 4: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude spanned by estimates in Thomas et al. (2016), Nadezhdina et al. (2016) and Park et al. (2024). Other observables in this plot are free to vary. In the 2D Probability Density Functions (PDFs), black (yellow) color indicates low (high) probability. The observed multiple peaks in the PDFs of $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, and phase lags reflect the different regimes of dissipation consistent with the observations (see Figure 2), resulting in a nonlinear forward model. Phase lags of $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ are in degrees, amplitude of the libration is in meters, and $J_{3}$ is divided by 106.
6.1.1 Plausible Interior Parameters and Observables
@@
Consistent with Section 3, the patterns of the posterior probability density functions (PDFs) of the amplitude and phase lag of $k_{2}$ largely differ from the PDFs of the amplitudes and phases of $h_{2}$ and $l_{2}$ (Figure 4). PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ have several peaks associated with tidal activity dominated in the core versus in the shell. The observed peaks in this figure are associated with the ranges of the plausible dissipation mechanisms both in the shell and in the core that are depicted in Figure 1. A different distribution pattern and peaks associated with different regimes of tidal activity are observed for $h_{2}$ and $l_{2}$, and their phase lags. This different pattern arises from the fact that $|h_{2}|$ and $|l_{2}|$ are largely insensitive to tidal activity in the core. The posterior PDFs show that while the $|h_{2}|$ and $|l_{2}|$ and their phase lags (Figure 4) are strongly correlated with each other (dominated by the response of the shell), their correlation with the amplitude and phase lag of $k_{2}$ depends on where the tidal dissipation dominantly occurs.
@@
We also obtain the ranges of the interior model parameters and the correlation between them associated with the observables (Appendix E, Figure 18). Previously, we ruled out mean crustal thicknesses around 5–10 km for Enceladus, as they would result in amplitudes of libration that are too large (see Section 4). Ice shell thicknesses $\approx$ 1–5 km would be compatible with the estimates of tidal heating and libration only if the mean effective viscosity of the shell is unrealistically low. Therefore, we conclude that the mean thickness of Enceladus’s shell is larger than 10 km. For completeness, we include very low effective viscosities for the shell and the core, which imply deformation in the viscous-dominated regimes. As a result, the PDFs of the viscosities of the shell and core indicate multiple peaks representative of tidal dissipation dominated by viscous deformation as well as shear (anelastic) relaxation (See also Figure 1). In the remaining analysis below, we consider viscosities of the ice shell and the rocky core that are close to or greater than their nominal relaxation viscosities ($\eta_{\rm shell}~>~10^{12}$ Pa.s and $\eta_{\rm core}~>~10^{11}$ Pa.s), i.e., the ice shell and the rocky core deform in the anelasticity-dominated deformation regime and not in the viscous-dominated deformation regime.
@@
We focus on the analysis of discriminating between two sets of geodetic observables and interior parameters that are associated with each of the two scenarios (Figure 5). In this figure, and in the remainder of this paper, we distinguish between core-dominated and shell-dominated tidal dissipation scenarios by associating them with the cases where at least 90% of the total heat is produced in the shell or the core. We compute the tidal heating in the shell using Equation (14). The PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ for the two tidal dissipation scenarios largely overlap, whereas the PDFs of $|h_{2}|$, $\epsilon_{h_{2}}$, $|l_{2}|$, and $\epsilon_{l_{2}}$ in the case of shell-dominated tidal dissipation differ significantly from those in the case of core-dominated tidal dissipation (Figure 5). In an inverse problem context, the degeneracy caused by the two different tidal dissipation scenarios reflected in the tidal Love numbers, as discussed earlier, leads to the non-uniqueness of the plausible interpretation of $k_{2}$ and $\epsilon_{k_{2}}$.
@@
By discriminating between the two tidal dissipation scenarios, we obtain a tight constraint on the viscosity of the shell in the cases where the tidal dissipation is dominated in the shell. Similarly, we constrain the viscosity of the core in the case where we assume that tidal dissipation occurs predominantly in the core (Figure 6). Based on the set of observables used in Figures 4 and 5, we find that the thickness of the ice shell, mostly constrained by the libration amplitude, is in the range of 15–40 km. Because we have conservatively chosen broad ranges for observables in this analysis, we have not obtained tight constraints on most of the other model parameters in this analysis. The PDFs and correlations between the interior structure parameters are presented in Appendix E.
@@
Figure 5: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux and the libration amplitude for tidal dissipation scenarios in the shell and in the core. The heat flux is assumed as $\rm 15~GW<\dot{E}<40~GW$, covering the estimated ranges in Hemingway et al. (2018); Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). In the 2D PDFs, the pink-gray and blue-yellow themes refer to dissipation in the shell and and in the core, respectively. All the variables are shown in logarithmic axes. Unit conventions are the same as in Figure 4.
@@
Figure 6: PDFs of ice shell and core viscosities for each of the two cases of tidal dissipation occurring dominantly in the shell (panel a) and in the core (panel b), corresponding to two sets of observables in Figure 5
We compare the PDFs of $h_{2}$ and $l_{2}$ and their phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) associated with the two heat generation scenarios to compute the minimum required precision of measurements to discriminate between the two dissipation scenarios (Figure 5). We compute the minimum measurement precision of $|h_{2}|$ and $|l_{2}|$ by calculating the separation between the two intervals determined by the 2-$\sigma$ uncertainties associated with the PDFs. Therefore, if we measure $h_{2}$ with a precision equal to the separation between the 2-$\sigma$ confidence intervals (or 1-$\sigma$, not shown in the figure) of the two PDFs, we obtain a confidence $\sim$95% (or 68% if the 1-$\sigma$ intervals are used) to identify the layer that dominates the dissipation budget. We present more detailed comparison of $|h_{2}|$ and $|l_{2}|$ and their phase lags in Appendix E.
@@
Due to the overlap between the PDFs of $|h_{2}|$ and $|l_{2}|$ associated with the two tidal dissipation scenarios using this set of observations, distinguishing between the heating in the shell or in the core is challenging. This large overlap is partly due to the large uncertainties in the observables considered in this case, and taking into account only the currently available constraints. The analysis implies that measurement of the maximum vertical tidal deformations with a maximum uncertainty of 62 cm can distinguish between the dominant region of tidal activity within the 1-$\sigma$ confidence level. Measurement of radial deformations with a maximum uncertainty of 56 cm can indicates the “more likely” region of heat generation, but it cannot definitively discriminate between the two scenarios. Alternatively, measurement of the tangential displacements within an uncertainty of 28 cm is required for 1-$\sigma$ confidence. The phase lags associated with the radial deformations ($\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$) indicate a better separation between the PDFs associated with the two regimes of tidal activity compared to the deformations and $|h_{2}|$. We find that measuring the phase lag angles of the tidal deformations with an accuracy of $1$ is required for 1-$\sigma$ confidence, and 4 is required for 2-$\sigma$ confidence in determining the dominant region of tidal dissipation (see Appendix E for more details).
6.1.2 Measurement Requirements Based on Park et al. (2024)
@@
Next, we focus on a narrower range of observational constraints by exclusively using the ranges of $\gamma$, $J_{3}$, moment of inertia (inferred from $C_{20}$ and $C_{22}$), total rate of heat loss (inferred from conductive shell thickness), and the density based on the analysis by Park et al. (2024) (see Table 1). We use 3-$\sigma$ uncertainties on all the observables. We present the correlation between the observables obtained based on this analysis and incorporating both tidal dissipation scenarios in Figures 7. Applying the narrower observational constraints based on Park et al. (2024) shows that $|k_{2}|$ and $\epsilon_{k_{2}}$ indicate a normal distribution with a single peak that encompasses both scenarios of tidal dissipation, making them indistinguishable based only on measurement of $|k_{2}|$ and $\epsilon_{k_{2}}$. However, $|h_{2}|$, $|l_{2}|$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ show two peaks associated with the two tidal dissipation scenarios. By separating the two tidal dissipation regimes (See Appendix E,  Figures 18b and 18c), we also quantify the preferred values of the observables that present the highest likelihood for each of the two heat generation scenarios based on the analysis of Park et al. (2024) (Table 3). We use these synthetic observations to perform inversions using plausible uncertainties expected from future measurements. We note that satisfying all constraints provided in Park et al. (2024) is possible when using 3-$\sigma$ uncertainties on the observables. Specifically, we find that the mean value of $J_{3}$ can be matched with the libration amplitude when considering 3-$\sigma$ uncertainty in the measured $J_{3}$. Our results can be compared to those presented by Genova et al. (2024) and Ermakov et al. (2021), where only tidal heating in the shell is considered. Also, Genova et al. (2024) use only the gravity measurements without exploiting the libration measurement as a constraint, which is expected to somewhat alter the results.
@@
Table 3: Favored values and uncertainties (3-$\sigma$) of observables calculated using MCMC sampling based on the available inferred observables given by Park et al. (2024). Also, expected uncertainties from future explorations are provided (see Section 6.4) which are used in the inversion in Section 6.2. The values in the parentheses in the expected measurements corresponds to a scenario with higher accuracy, as discussed in the text. Expected uncertainties are associated with 2-$\sigma$ error. Note that if we do not separate the two dissipation regimes, the $|k_{2}|$ and $|\epsilon_{k_{2}}|$ have single peaks at 0.027 and 34, respectively.
@@
$\mu_{\rm shell}$ GPa 242-4 242-4
$\rho_{\rm shell}$ kg/m3 700950700-950 700950700-950
$\rho_{\rm ocean}$ kg/m3 9201250920-1250 9201250920-1250
$R_{\rm core}$ km 183213183-213 183213183-213
$\mu_{\rm core}$ GPa 4704-70 5805-80
$\rho_{\rm core}$ kg/m3 216024002160-2400 216024002160-2400